SITE 1094
(Proposed Site TSO-7C)

The voyage to Site 1094 north of Bouvet Island began with the wind gusting to 38 kt and seas and swell nearly at right angles to the ship's track. Rolling heavily, the vessel made only about 6 kt on her southerly course for the remainder of the evening. At 2330 hr, it was necessary to slow the shafts and to alter course because of excessive vessel motion.

At 0030 hr on 20 January, the drilling crew reported that the forward retaining pin and the two safety slings of the lower guide horn (LGH) in the moonpool had broken and that the port half of the LGH was swinging free about its hinges on the aft pin. The ship was slowed to ~2 kt in the heavy weather while attempts were made to restrain the LGH. These attempts were unsuccessful, and speed was increased to 5 kt in the direction of the drill site, which was still about 150 nmi away. At 0315 hr, the aft pin failed just above the water line and allowed the upper part of the unrestrained section of the LGH to move forward and outboard until it came to rest against the wall of the moonpool. The port LGH section (of undetermined length, later determined to be complete) remained attached to the starboard half at a lower corner by a section of the aft retaining pin. The displaced portion was securely wedged across the moonpool and remained in place as the ship proceeded at full speed toward the drill site in rapidly improving weather.

Visual inspection indicated that the center of the moonpool was unobstructed and that drilling operations were feasible to the extent allowed by the 20-ft reduction of the port LGH. Operations thus continued according to plan, and a positioning beacon was launched at 2000 hr on 20 January.

Hole 1094A
While the vessel was stabilizing its position on the beacon, a stand of drill pipe was picked up and run through the moonpool to confirm that the center well was unobstructed. The signal from the positioning beacon was weak and considered unreliable, so the backup beacon was launched at 2050 hr while the BHA was being assembled.

Weather and swell conditions improved dramatically to nearly flat calm as spud time approached, allaying concerns about stricter operating limits on vessel motion that were expected from ODP engineers. The pipe trip was slowed somewhat because stands of drill pipe had shifted on the piperack during the rough weather and had to be repositioned for handling by the automatic racker.

At 0408 hr on 21 January, Hole 1094A was spudded with an APC core from 2814 mbrf. The initial 4.6-m core set the seafloor depth at 2818.9 mbrf. APCT temperature readings were taken with Cores 1H (for bottom water temperature), 4H, 6H, 8H, 10H, 12H, 14H, and 16H. The majority of cores were oriented azimuthally as the one remaining operational instrument permitted. Coring results were affected adversely by the accumulation of ice-rafted debris in the hole and by persistent core liner failures. (Nearly 100% of the liners failed in some manner.)

Core 18H indicated incomplete stroke and produced no recovery. A 3-m interval was drilled in case the refusal was due to a thin hard stratum. Core 19H also had incomplete stroke but recovered 8 m of core which apparently was a stack of two sections resulting from two successive stabs with the corer. As APC refusal apparently had been reached, a Davis-Villinger temperature probe run was made to complete the temperature gradient profile.

One additional attempt was made to obtain an APC core before abandoning the hole. Several attempts to actuate the corer were unsuccessful and the APC was recovered with the shearpins intact. Coring in Hole 1094A was discontinued at the request of the scientific party and the drill string was withdrawn from the hole.

Hole 1094B
The rig was offset 10 m, and Hole 1094B was spudded with a seafloor APC core at 0015 hr on 22 January. Core recovery provided close agreement with the seafloor depth of Hole 1094A.

The anticipated guidelines for coring with a damaged LGH were received while Hole 1094B was being cored. Roll and pitch limits were reduced, as expected, and the use of knobby drilling joints was mandated for all drilling/coring operations. The knobbies were put into service beginning with Core 4H.

Cores 1H and 2H gave essentially full recovery, but recovery was poor in Cores 3H and 4H. All four cores gave pressure indications of incomplete stroke. Because of the good results in Hole 1094A, the presence of ice-rafted debris in the hole was blamed for the unsatisfactory performance in Hole 1094B. To make the best use of remaining operating time, the decision was made to pull clear of the seafloor, offset 10 m to the opposite side of Hole 1094A, and start over.

Hole 1094C
A third successful mudline core was collected with Core 1094C-1H, shot at 0443 hr from 2818 mbrf. All three seafloor depths agreed within 1 m. Problems with APC performance again were experienced as eight cores were attempted and six of them gave pressure indications of incomplete stroke. The two cores that indicated full stroke had no recovery. After Core 8H, which recovered 1.3 m of sediment, the drill struck hard material after reaming 7.2 m toward the next core point. No measurable penetration was made after 20 min of rotation. Again the coring problems were attributed to ice-rafted material in the sediment, with the final hard streak interpreted as a boulder. Coring attempts were abandoned, and the bit was raised above the seafloor.

Hole 1094D
The ship was offset about 20 m to the northwest of the coordinates of the earlier holes. The hole was drilled to 19.1 m below inferred seafloor depth before the APC was deployed. Continuous APC cores 1H through 13H were taken to 142.6 mbsf before it was necessary to replace the knobby joints with standard drill pipe. When the bit reached TD following the 2-hr short trip, about 1.5 m of fill were found in the hole. A mud flush was pumped concurrently with Core 14H, following the trip.

At the time Core 16H was being recovered from 171.1 mbsf (0215 hr on January 23) a partial load shedding to the ship's dynamic positioning (DP) system occurred, presumably the result of an unexpected drift-off (2% yellow alarm) in relatively calm seas and moderately windy weather (30-40 kt gusts). Within 3 min, position was recovered to 1%. The load shed in conjunction with exceeding operations limits (4° pitch or roll) led to the decision by the Operations Manager to terminate the scientific drilling operations of Leg 177. Coring at Site 1094 was therefore terminated about 1.5 days ahead of schedule.

When the drill string had cleared the seafloor, weather and motion conditions were essentially unchanged, and it was deemed safe to run the coring line back down the pipe so that it could be sprayed with protective coating while being retrieved. The drill string then was tripped during improving weather conditions, and the BHA was broken down and stored for transit. The bit arrived on deck at 1145 hr on 23 January.

An extensive effort then was made to secure the broken LGH-half in the moonpool prior to the transit to Punta Arenas. A retaining frame consisting of two 17-ft I-beams in an open cross configuration had been fabricated on the rig. The frame, complete with padeyes and heavy slings, was keelhauled over the side of the vessel (the moonpool doors could not be opened) and suspended from the main traveling block by cable slings through the center well of the moonpool. The block then was raised to engage the bottom of the LGH-half with the frame. After several attempts, the frame was determined to be supporting the LGH. Efforts to lift and realign the LGH-half were unsuccessful, but a tension of 30 kips on the slings, in addition to the attachment of the lower corner at the locking pin, seemed to stabilize the assembly and prevent motion relative to the moonpool wall and the starboard half of the LGH. That appeared to be the best possible preparation for the long, rough transit ahead, and the rig floor was secured with the driller at his station to monitor the tension on the support frame.

The vessel had remained in full DP mode for the securing operation, and an additional 3 hr were required to recover the positioning beacons, house the hydrophones, raise the thrusters, and do the protective maintenance for transit that normally is done during the final pipe trip. The vessel departed Site 1094 at 1845 hr on 23 January.

To 177 Transit to Punta Arenas

To 177 Table of Contents

ODP Publications

ODP Home